what is the minimum value of 3sin 6x + 3cos6x ??   

Grade:Select Grade

2 Answers

Lab Bhattacharjee
8 years ago
\text{Using AM, GM inequality,} \\ \dfrac{3^{\cos6x}+3^{\sin6x}}2\ge \sqrt{3^{\cos6x\cdot\sin6x}} \\ \text{Now,} \cos6x\cdot\sin6x=\dfrac{\sin12x}2 \\ \text{ and} \sin12x\le1
Lab Bhattacharjee
8 years ago
\text{There was typo in the previous answer} \\ \text{Using AM, GM inequality } \dfrac{3^{\cos6x}+3^{\sin6x}}2\ge\sqrt{3^{\cos6x+\sin6x}} \\ \text{ Now,}\cos6x+\sin6x =\sqrt2\cos(6x-45^\circ)\le\sqrt2